OpenStudy (anonymous):

Given that f(x) = 4x-3 and g(x) =2x-1/3, solve for g(f(2)). 1 3 9 15

OpenStudy (anonymous):

it's 15 i think

OpenStudy (anonymous):

\[f(x) = 4x-3 \text{ and } g(x) =\frac{2x-1}{3} \text{, so } f(2) = 5 \text{, so } g(f(2))=g(5)=3\]